Write out the products including the physical state and balance the following reactions by using the...

  1. Write out the products including the physical state and balance the following reactions by using the principles of metathesis reaction:

**PLEASE SHOW YOUR WORK I AM TRYING TO LEARN HOW TO DO THESE AND IT JUST ISN'T CLICKING**

      a)               Mg(C2H3O2)2(aq) + Cs3PO4(aq) ==>

      b)               FeCl3(aq) +   AgNO3(aq) ==>

      c)               Cr(NO3)3(aq) + Ba(OH)2(aq) ==>

      d)              Fe(OH)2(s) + HBr(aq) ==>

      e)              Ca(OH)2(aq) + H2SO3(aq) ==>

  1. ) HC2H3O2(aq) + LiOH(aq) ==>

      g)               (NH4)2CO3(aq) + Sr(OH)2(aq) ==>

h)               Mg(s) + Cu(NO3)2(aq) ® Mg(NO3)2(aq) + Cu(s) ==>

      i)                Ba(s) + HCl(aq) ® BaCl2(aq)   + H2(g) ==>
Answer #1

(A) 3 Mg(C2H3O2)2(aq) + 2 Cs3PO4(aq) = 6 CsC2H3O2(aq) + Mg3(PO4)2(aq)

(B) FeCl3(aq) + 3 AgNO3(aq) = 3 AgCl(s) + Fe(NO3)3(aq)

(C) 2 Cr(NO3)3 + 3 Ba(OH)2 = 2 Cr(OH)3 + 3 Ba(NO3)2

(D) Fe(OH)2(s) + 2 HBr (aq) = FeBr2 (aq) + 2 H2O

According to HOMEWORKLIB POLICY only 1 question with 4 sub part per post is allowed. please post the rest question seperately.

PLEASE LEAVE THE POSITIVE FEEDBACK IF YOU LIKE THE SOLUTION